OpenStudy (anonymous):

On a map, 90 miles is represented by 3 inches. A measure of inch on the map would represent _________miles.

jimthompson5910 (jim_thompson5910):

3 inches = 90 miles 3/3 inches = 90/3 miles 1 inch = 30 miles So one inch on the map represents 30 miles in real life